Assume the demand function for good X can be written as Qd = 80 - 3Px + 6Py + 10I, where Px = the price of X, Py is the price of Y and I is consumer income. If the price of Y decreases by 5 dollars, what is the change in Px have to be in order to keep the quantity demanded of X unchanged by the change in the price of Y?

A) decrease by 2.5 dollars B) decrease by 10 dollars C) increased by 10 dollars D) decrease by 1 dollar
Answer #1

Demand for X = Qd= 80-3Px+6Py+10I

Py falls by 5 dollars

New Py = (Py-5)

New Qd = 80-3Px+6(Py-5)+10I

=80-3Px+6Py-30+10I

50-3Px+6Py-10I

To keep quantity demanded unchanged price of X should fall by 10 dollars (hit and trial method)

New Px= Px-10

New Qd= 50-3(Px-10)+6Py+10I

= 50-3Px+30+6Py+10I

=80-3Px+6Py+10I

which is original quantity demanded.

So price of X should decrease by 10 dollars